OpenStudy (anonymous):

find the area of the circle pictured. Use a- 3.14r^2

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

(2n-3)(2n-3)*(3,14)=(4n^2 -2*2n*3 +9)*(3,14)=

OpenStudy (anonymous):

= what?

OpenStudy (anonymous):

Is that a 2n-3 for radius?

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

a. a=12.56n^2-27.68n+28.26 b. a-12.56n^2+28.26 c. a= 6.28n^2-9.24

OpenStudy (anonymous):

\[3.14(2n-3)^2\] Do you have to expand?

OpenStudy (anonymous):

Apparently you do.

OpenStudy (anonymous):

it a got it